the question is: if we begin a graph of f(x)=the square root of x, shift 4 units to the right, shrink vertically by a factor of 1/2 , and shift upward 10 units, write an equation for the transformed graph.
so far I have:
(the square root of x)- 4(four units to the right) + 10( 10 units upwards)
I do not know how to do with the part that says shrink vertically by 1/2
Answer #1
The question is: if we begin a graph of f(x)=the square root of x, y = sqrt(x) -------------------- shift 4 units to the right, y = sqrt(x-4) -------------------- shrink vertically by a factor of 1/2 y = (1/2)sqrt(x-4) ------------------------------------ , and shift upward 10 units, y = (1/2)sqrt(x-4)+ 10 ================================== Cheers, Stan H. --------------------------------
answered by: saud